How to Obtain an Academic Position (Or at Least, How to Try to Obtain One)

Larry Hubert/Carol Nickerson (with a little help from our friends in the

Quantitative Division)

Where to find ads for position openings

* Major sources

— American Psychological Association (APA) primarily ads for academic, research, and post doc positions (plus clinical openings) printed ads in each month’s issue of the APA Monitor online ads at psyccareers.apa.org/ online ads are searchable by geographic location and category; see especially the category “Quantitative/ Mathematical/ Psychometrics/ Statistics”

— American Psychological Society (APS) primarily ads for academic, research, and post doc positions (fewer clinical openings) printed ads in each month’s issue of the APS Observer (I believe this may be “the best” general place for all of academic psychology.) online ads at www.psychologicalscience.org/jobs online ads are searchable by keyword

— Chronicle of Higher Education (CHE) ads for both academic and non-academic positions in all fields (not just psychology) printed ads in each week’s issue of the Chronicle of Higher Education online ads at chronicle.com/jobs/ online ads are searchable by position type, discipline/field, state/region, institution, and keyword

— American Statistical Association (ASA) ads for both academic and non-academic positions printed ads in each month’s issue of the AMSTAT NEWS the small (65-word), but not the large, ads from AMSTAT NEWS are also available online at www.amstat.org/ads online ads are searchable by geographical location and keyword

other online ads (not appearing in AMSTAT NEWS) are available on the ASA “JobWeb” at

jobs.amstat.org/search/results

online ads are searchable by job category, job level, type of job, geographical location, and keyword

* Special interest groups

— Society for Judgment and Decision Making (SJDM)

printed ads are in the quarterly newsletter; other ads in the mailing list and the mailing-list archives; all available online at

www.sjdm.org — Psychometric Society online ads available at www.psychometrika.org/members/PMjobs.html — National Council on Measurement in Education (NCME) online ads available at www.ncme.org/careers/opportunities.cfm

— American Educational Research Association (AERA) printed ads in each month’s issue of the Educational Researcher online ads available at www.jobtarget.com/home/index.cfm?site id=557 * Other sources

the Quantitative Division faculty regularly send ads that they receive to the “quantoids” mailing list [handout example]

the bulletin board between Carolyn’s (429) and Larry’s (431) offices, and the bulletin board near the conference room (318) on the third floor

- the Psychology Department mailing list

primarily ads for academic positions and post docs in all areas of psychology

from IMP/Horde Webmail (cyrus.psych.uiuc.edu), open folder “Public/Psych-Notes/Job-opportunities”

Six components of the application for an academic position

(1) cover letter

- can be brief if (2), (3), and (4) below are comprehensive and well-written

- might tailor cover letter to a particular position or set of positions (e.g., positions in educational psychology versus positions in quantitative psychology)

- indicate that you have requested that reference letters be sent (and from whom) if the position ad requires them to be sent

(2) curriculum vitae (CV or vita)

- a content outline and lots of good tips are available at

www.psychwiki.com/wiki/Curriculum Vitae [handout] - look at CVs written by other persons for content and layout ideas

- read the “CV Doctor” advice articles on the Chronicle of Higher Education website chronicle.com/jobs/tools/cvdoctor/2007/

- ask a faculty member to review your CV (3) statement of research interests

- as individual as you are

- helpful hints at www.vpul.upenn.edu/careerservices/gradstud/ Writing Research Statement.ppt [handout] (4) statement of teaching interests and/or philosophy

- helpful hints at

gradpsych.apags.org/nov05/teach.html [handout]

You can also obtain a formal teaching “certificate” within the Psychology Department, along with your own TA responsibilities (by working with Sandy Goss-Lucas). This would look good on your teaching statement as a measure of your commitment to instruction.

Graduate Teaching Certificate

www.oir.uiuc.edu/Did/TAs/GTC.htm

Advanced Graduate Teaching Certificate

www.ori.uiuc.edu/Did/TAs/AGTC.htm (5) copies of publications — reprints of published articles or book chapters, preprints of “in press” articles or chapters, copies of submitted manuscripts

- usually three are sufficient

- print double-sided pages to save postage (6) a commitment from three persons to serve as a reference

- usually one of these will be the faculty member supervising your dissertation [handout]

- other persons might be faculty members on your thesis committee and/or from the Quantitative Division (e.g., someone for whom you have served as a teaching assistant), other divisions, or the Statistics Department; or a summer internship mentor

- Provide each of these persons with a list of positions to which you are applying, noting the type of position, the person to whom the reference letter should be addressed, and the application submission deadline (and perhaps a copy of the position ad also). Also, provide a copy of (2), (3), and (4) as an aid to writing the reference letter.

- Some position ads will require that you request that reference letters be sent at the same time that you submit your application. Others will require only the names and addresses of the persons writing your reference letters. Reference information should be included in your cover letter and near or at the end of your CV.

- Give the person writing you a reference letter adequate lead time, especially for the first letter.

- Some faculty are reluctant to write reference letters for more than one applicant for a particular position. You may need to co-ordinate a “reference strategy” with the other job applicants and the faculty from the Quantitative Division.

* examples of (1), (2), (3), and (4) from former Quantitative Division graduate students and post docs [handouts] * The Psychology Department will reimburse graduate students (but not post docs) for up to $150 in job-application mailing expenses; see the memo from David Irwin for the procedure to follow. [handout] Timeline

* Position ads begin to appear in August, continuing through the fall, with most application deadlines in the fall, some as early as 15 October, although some may be as late as February or March.

* Pay attention to application submission deadlines; these are crucial.

* Some departments may start interviewing before the application submission deadline so don’t wait until the very last minute to submit your application.

* Interviews usually begin in the late fall, with a break between 15 December and 15 January, and then continue through February and perhaps March.

* Thus, to consider entering the job market, you should have defended your dissertation proposal by the middle of the fall semester.

* Ads appearing after December will usually be for post docs; industry, government, or laboratory positions; temporary (rather than tenuretrack) positions (e.g., a replacement for a faculty member going on sabbatical); and so on.

* The status of a particular position opening may be posted on the “wiki”

psychjobs.wikidot.com/psych-job-search-07-08

Thanks to Greg Webster for this tip!

The job talk

* If you receive a request for an on-site job interview, you will be asked to give a research “job talk” and perhaps another talk, probably either a “teaching talk” for a teaching-oriented position on either a statistical topic or a substantive topic, or a specialized “quant talk” (“two theorems and a couple of lemmas”) if the Psychology Department to which you are applying has a separate Quantitative Division. If you are clinical, you will almost always need to give a “clinic talk” (usually on a client(s) that you are seeing).

* The primary job talk (usually) should be geared for a general audience (not too many equations) and made interesting with examples, possible applications, etc.

* The job talk has a disproportionate influence on the hiring decision. In fact, it is ridiculous how much emphasis is placed on the job talk — why does no one remember the unreliability of single behavioral instances?

Do the very best you can. Practice your job talk before the interview by giving it at the brownbag seminar or by rounding up your fellow graduate students, post docs, and faculty to listen. Inviting some persons from outside the Quantitative Division is a good idea. Ask the audience for honest feedback.

* Job talks are usually scheduled for 45 to 50 minutes, with 10 to 15 minutes for questions.

* Make arrangements beforehand with the hosting department for any audio/visual equipment that you might need. [laptop to borrow, for instance] The site interview

* Before an invitation for a site interview is extended, you may receive a request for a preliminary telephone or “screening” interview, which is used to whittle a “short list” of possible candidates down to the three or four invited for a site interview.

* A site interview may be for 1, 1 1/2, or 2 days. The schedule may look something like this:

Arrive Sunday: met at the airport by the chair or a member of the search committee

Monday/Tuesday: “meet and greet” faculty and graduate students, perhaps tour the facilities Monday afternoon: the job talk

Monday evening: dinner with selected faculty

Tuesday morning: possible teaching, clinic, or quant talk

Tuesday noon: lunch with graduate students

Tuesday afternoon: exit interview with department head/chair

Tuesday evening: departure

* During the exit interview, you and the department head/chair will likely beat around the bush with respect to salary and start-up arrangements. Good research-oriented universities are currently offering starting salaries of about $60,000 to $65,000. The APA and the ASA publish annual salary surveys. research.apa.org/salsurv.html

www.amstat.org/profession/ index.cfm?fuseaction=main

* If you are part of a dual-career couple, the exit interview is the place to mention it, but don’t over-emphasize it. Explain what your spouse/partner does and the type of position in which he or she is interested.

Travel arrangements for site interviews

* Usually you will make your own transportation arrangements and then be reimbursed by the hosting department. You may or may not need to make hotel arrangements as well, but you probably will not need to make meal arrangements.

* NEVER agree to paying the expenses of your interview yourself. You are better than that!

Also, you shouldn’t do “if you take our offer, we will pay for your trip”. Or “if we don’t make an offer, we will as well”. But if “we make an offer and you turn us down, it is on your nickel.”

The APA advertising policy stipulates that position opening ads must indicate whether interview expenses are not to be fully paid by the interviewer.

* Get a credit card if you don’t have one. Sometimes there is a lengthy delay between the interview trip and reimbursement.

What to talk about when you “meet and greet” faculty and graduate students during the site interview

* Don’t forget that you are on an interview, even during meals, casual conversations, parties, and graduate student meetings. Don’t drink alcohol during the meals; don’t say inappropriate things (e.g., sexually charged, political, ethically insensitive, and so on), even to graduate students.

* Know at least a little about the research areas of the faculty that you will be meeting, especially those working in your area.

* Invariably you will be asked if ”you have any questions”. Some possibilities to fill the void:

- “How is the department structured? How big is it? How many faculty, graduate students, undergraduate students?”

- “What is the usual teaching load?” (and to the chair of the search committee, perhaps: ”What courses might you expect me to take over?”)

- “What is the climate of the department?”

- “What are the undergraduate and graduate students here like?”

- “What is it like to be a woman faculty member here?” (probably best asked by a woman candidate of a woman faculty member)

- “What proportion of assistant professors make tenure?” “What are the requirements for tenure?”

- “What is the town like?” (“It is not the end of the world but you can see it clearly from here.”)

- If talking to a fellow quantitative psychologist, inquire tactfully whether the department will be supportive of your research program or whether the position is primarily one of (teaching/consulting) service to the department — this latter expectation is commonly called being the “mule.”

It’s OK to ask the same question more than once; different people may offer different perspectives.

* There are some questions that legally you cannot be asked during a job interview, including marital status, citizenship, visa status, pregnancy plans, etc. See the “UIUC Guidelines for Pre-Employment Inquiries”. [handout] The job offer

* If you receive a job offer, it will usually arrive in two parts:

(1) a letter from the Dean formally offering you the position, indicating a salary and reminding you not to get too comfortable until you obtain tenure. (Per the agreements all institutions have with the AAUP, you cannot remain as an assistant professor for more than seven years. The tenure decision is usually made in the sixth year. You can usually remain a terminal seventh year while searching for a new position.)

(2) a letter from the department head/chair specifying other details, such as start-up research funds, teaching load, office and lab arrangements, etc.

See as an example the letter that I (when Acting Department Head) sent to an incoming faculty member. [handout] Hope that you are not asked for a specific and detailed list of what you might need in the way of start-up, and instead are just given a lump sum from which to buy your start-up needs.

If you are asked for a list:

— computer(s) for lab/office/home; laptops; printers; other necessary hardware

— software, depending on what is site-licensed there; MATLAB, SYSTAT, SPSS, SAS, AMOS, Office Pro, and so on

— subject monies

— some conference travel funds

— supply and expenses budget for photocopying, postage, miscellany supplies

Try to arrange a deal in which the start-up fund doesn’t disappear after a year or two; this may be very hard to do, however * Typically, you will have about two weeks to make a decision. Sometimes it is possible to request an extension to this time limit if you are anticipating other job offers. But other times, they may play hardball.

* Generally, it is not a good idea to pit one or more offers against each other. Even if you win and gain an incrementally better arrangement than the other assistant professors have, everyone will hate you coming in — not a very good beginning environment for you.

* NEVER EVER EVER EVER accept a job offer and then turn it down for another (better) offer. You will be disowned by your field; if we could do it, you would be placed out on a rock to wither away.

Miscellaneous tips

* Attend job talks in the UIUC Psychology Department, even if they are not in your research area. See how it’s done (and maybe how it’s not done).

[handout]

* Begin keeping a “professional development” folder during your first year as a graduate student. This will make preparing your CV and teaching and research statements much easier. In this folder, keep records of

- research assistantships

- other research projects

- teaching assistantships and course evaluations - published articles

- journals for which you have written reviews (and the reviews)

- conference presentations; brownbag presentations

- transcripts (for a record of courses taken that you may be able to teach someday)

- honors, awards, scholarships, fellowships, “Incomplete List of Teachers Ranked as Excellent by Their Students”, etc.

- little notes of appreciation that you might receive (nice to re-read these on a “down” day)

- memberships in professional organizations - service activities

- statistical consulting gigs

- and whatever ...

Begin your CV during your first year of graduate school and update it regularly.

* Publish, publish, publish – with your research mentor, by yourself, with other graduate students (the latter is really impressive!). Read “How your publication record is evaluated?” available at

www.psychwiki.com/wiki/

How your publication record is evaluated

* Submit a presentation proposal to the Bohrer workshop hosted by the Statistics Department during the spring semester. Everyone gets an award. (“At last the Dodo said, ‘EVERY-BODY has won, and all must have prizes’.”)

* Learn Tex/LaTex and MATLAB.

* Dual-career couples will have a more difficult time finding academic positions; the higher the academic caliber of both persons, the easier the job search will be.

Dual-career couples generally may have more opportunities in large cities that have several colleges and universities in the environs (e.g., New York, Boston).

Illinois, for example, has a very good dualcareer program, with partial finding (1/3 of the line) from the Provost’s office. I know of five dual-career hires in the last three years, just in Psychology (because I was responsible for pushing these through); these are balanced across sex (i.e., who is the leading spouse and who is the trailing spouse). Other institutions may not be so enlightened — Illinois has to be, given its location.

A “portable” spouse or partner (i.e., someone who is employable just about anywhere, usually not an academic) is an asset.

* The best advice that I have ever received (from Frank Baker, when I started my career at the University of Wisconsin):

*** YOU MUST SEPARATE YOUR PRO-FESSIONAL WORTH FROM YOUR PER-SONAL WORTH ***

Believe me, the tenure process and the review/publication process are brutal (hear me, just brutal), and can be devastating if you don’t maintain this separation.

* Useful books Darley, J. M., Zanna, M. P., & Roediger, H. L. III. (2003). The compleat academic: A career guide (2nd ed.). Washington, DC: American Psychological Association. Heiberger, M. M., & Vick, J. M. (2001). The academic job search handbook (3rd ed.). Philadelphia, PA: University of Pennsylvania Press.

* Useful websites

- the APA career website at psyccareers.apa.org/

- the Chronicle of Higher Education career website at chronicle.com/jobs/ - the UIUC Graduate College Career Services Office website at

www.grad.uiuc.edu/careerservices